Operations Report – Nov 9th

[title Operations Report – November 9th]

[category operations-report]

Crew 214 Operations Report 9-11-2019
SOL: 13
Name of person filing report: Andrew Wheeler
Non-nominal systems: Opportunity, Suit 10
Notes on non-nominal systems: Opportunity brakes do not work, Suit 10 is not functioning
Generator: run
Hours run: 13h
From what time last night: 1815
To what time this morning: 0715
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night): 70%
Diesel Reading – 55%
Station Propane Reading – 76%
Ethanol Free Gasoline: N/A
Water (loft tank): 41 gallons
Water Meter: 1456896 units
Water (static tank): 150 gallons (not filled)
Static to Loft Pump used – yes
Water in Green Hab: 225 gallons
Water in Science Dome: 0 gallons
Toilet tank emptied: yes
Deimos rover used: not at hab
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Spirit rover used: no
Hours: 109.2
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 100%
Currently charging: yes
Opportunity rover used: no
Hours: 65.6
Beginning charge: 100%
Ending charge: 100%
Currently charging: no
Curiosity rover used: no
Hours: 115.0
Beginning charge: (Before EVA) : 100%
Ending charge: (On return from EVA, before charging): 100%
Currently charging: yes
Notes on rovers: no
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used: 0 Gals
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director
HabCar used and why, where? N/A
CrewCar used and why, where? CrewCar to Grand Junction and return for Crew 215
General notes and comments: N/A
Summary of internet: Functional
Summary of suits and radios: Nominal
Summary of Hab operations: All systems nominal
Summary of GreenHab operations: Atmospheric equalization to the green hab was carried out during the day, plants were watered. The heater automatically came on at 5pm.
Summary of Science Dome operations: All systems nominal.
Summary of RAM operations: Not used
Summary of any observatory issues: Nominal
Summary of health and safety issues: All crew healthy
Questions, concerns and requests to Mission Support: Static tank not filled. Water heater unusable. Opportunity is currently unusable.

Copyright © The Mars Society. All rights reserved. | Main Site